Activity at the volcano was at about the same level same as yesterday with dust levels low and rockfall events dominant.
Between 16:00 yesterday and 16:00 today there were 14 hybrid earthquakes, 3 long period earthquakes, 2 volcano-tectonic earthquake and 38 distinct rockfall signals. Eight of the long period earthquakes preceded rockfalls. There were two periods of elevated seismic activity during which most of the rockfalls occurred. These were experienced between 6:30 pm and 8 pm last night and between 6 am to 8 am this morning. Two regional earthquakes were recorded by the seismic network at 1:33 am and 4:25 pm local time. The location of these events have not yet been determined.
The periods of heightened seismic activity were again correlated with vigorous ash venting and generation of steam plumes. Overcast conditions prevented any clear views of the dome.
Occupation of GPS stations by the MVO and University of Puerto Rico at Harris, Lookout, St. George's Hill, Perche's Mountain and Lee's Yard ended today. The results of these measurements are currently being processed.
A new electronic tilt station was installed at the top of Gages Mountain today. This station along with the one recently installed at Hermitage would improve the MVO's ability to monitor deformation of the northern flank of the volcano.
Dust levels at all of the monitoring sites remained low today.
Dr. Gill Norton of the British Geological Survey left today after another three month stint as deputy Chief Scientist at the MVO.